OpenStudy (marcelie):

for the following exercises find the x and y intercepts for the functions ... f(x)= x/x^2-x

OpenStudy (campbell_st):

well there are no x intercepts... looking at the denominator \[x \neq 0, 1\] because \[f(x) = \frac{x}{x^2 -x} = \frac{x}{x(x -1)}\] x = 0 is a point of discontinuity and x = 1 is a vertical asymptote. when you simplify you get \[f(x) = \frac{1}{x - 1}\] this has a horizontal asymptote at y = 0 so looking at the y- intercept let x = 0 \[y = \frac{1}{0 - 1}\] just simplify it. hope it helps

OpenStudy (marcelie):

why isnt there no x int

OpenStudy (marcelie):

oh nvm i got it lol thank you !!!